In a shipment of 10 laptops, 3 are defective. If 3 laptops from this shipment are randomly selected and not replaced, what is the probability that all 3 are defective? Record your answer as a fraction in simplest form.


1
Expert's answer
2021-12-12T17:44:39-0500
"\dfrac{3}{10}(\dfrac{2}{9})(\dfrac{1}{8})=\dfrac{1}{120}"